OpenStudy (anonymous):

The length of a rectangle is 4 centimeters less than twice its width. The perimeter of the rectangle is 34 cm. What are the dimensions of the rectangle?

OpenStudy (kash_thesmartguy):

Here, we need to solve it by making formulas. Any idea what any of them can be?

OpenStudy (anonymous):

nope, all I have are the choices

OpenStudy (kash_thesmartguy):

Where are the choices? Please post full questions from here onwards

OpenStudy (kash_thesmartguy):

As @pooja195 said: 1st formula: 2w-4=l

pooja195 (pooja195):

x - width 2x-4 - length 2x+2(2x-4)-34 2x+4x-8=34 6x=42 x=7 2*7-4=10

pooja195 (pooja195):

Dimensions of the rectangle are 7 and 10 x- width and x=7 so 7 is width